0850: Day starts: Drink 1L of water to hydrate body and counteract dehydration caused by long term exposure to dry, cool air conditioned laboratory.
0900: Confer with colleague. Discussed recent results with a colleague and outline strategy for samples to grow today.
0915: System checks. Inspected the ion gauge pressures, the status of all pumps and LN2 level. Inserted MIG into beam path in preparation for BEP flux measurements.
0916: Outgas Ga. Executed a program in the Ga cell’s PID controller than ramps up to 975°C in 30 minutes, holds for 30 minutes and then ramps down to 950°C and holds.
0920: As flux check. As shutter and valve were opened. Fluxes were taken at several valve positions consistent with previous days. Fluxes are recorded in the “flux check record” tab of the log spread sheet.
0946: As flux check completed. Gathered flux values were compared with legacy data for continuity checks. As shutter is closed and background As4 is pumped from the system. Vacuum recovery progresses prior to Ga flux check.
1000: Ga flux check. Ga shutter was opened for 60 seconds and the flux was recorded, averaged over 10 seconds. The shutter was then closed, the background flux was sampled, subtracted and the “flux check record” tab of the log spread sheet was updated. Current flux value is compared to legacy data for continuity.
1010: Sample 1: Start. Transferred sample from outgas stage to growth chamber and manually began recipe. Notice that yesterday’s “development” of the control software has resulted in the substrate rotation behaving very strangely! Disabled lift motor just in case. Contacted technician.
1020: Sample 2 Outgas. Transferred sample 2 from FEL to outgas stage and executed outgas program in PID controller.
1230: Sample 1: Complete. Transferred sample from growth chamber to FEL.
1240: Lunch and consumption of another litre of water.
1245: Technician investigates substrate rotation issue. The technician investigated the virtual interface communication and investigated the rotation issue. The fault was not found. The rotation will be operated manually (with the virtual interface deactiveated).
1515: Sample 2: Start. Transferred sample from outgas stage to growth chamber and manually began recipe.
1600: Platen 5. Received platen 5 back from etching. Loaded a Ta plate in preparation for outgas.
1610: Platen cleaning with P50 dry abrasive sheet. Most of material removed whilst plate was dry. WOrk was performed in a fume cupboard and a face mask was needed as there was much dust. Clean ok.
1655: Sample 2: Complete. Transferred sample from growth chamber to FEL.
1700: Technician investigates substrate rotation issue (2). The technician continued to investigate the substrate rotation issue and rectified the communications problem. The virtual interface can be utilised once more.
1730: Disperse NWs for TEM. As outlined below.
- Cleave a small piece of wafer
- Place 5ml of IPA in a lidded container
- Insert piece of wafer
- Scratch the surface with tweezers to snap many NWs
- Place small glass bowl in ultra sonic bath
- Add small amount of water to the bowl
- Carefully place the lidded container in the bowl
- Ultra sonic for 5 minutes, then…
- Place a TEM grid on a piece of filter paper
- Check TEM grid is the correct way up in the microscope
- Use a pipette to remove a small amount of fluid from the lidded container
- Carefully wet the TEM grid from the solution in the pipette
- Place filter paper and TEM grid on the hot plate
- Repeat wetting 5 – 8 times, leaving to try between each step
- Remove filter paper and TEM grid from hot plate
- Using vacuum tweezers (carefully) place TEM grid on a glass slide
- Check NWs are on the TEM grid with optical microscope
- Place TEM grid in a small box (blue lid)
- Store in desiccator
1830 Sample load/unload. The two samples in the FEL (grown today) were removed and placed in labelled boxes. Two new Si(111) substrates were mounted along with Platen 5 (fitted with a Ta disc) and the FEL pumped down manually and the degas cycle started . The intention to implement “automatic FEL pump down and degas” was reaffirmed whilst shaking a fist at the heavens. Samples were passed to a colleague for SEM analysis.
1900: Update log. Sample locations were updated on the “locations” tab of the log spread sheet to reflect new sample locations and status.
1910: Day ends. Drink another litre of water. Go home.
Total 10.33 hours. Growth: 2 samples. Samples/hour: 0.19